OpenStudy (anonymous):

How many ounces are in 6 pounds?

jimthompson5910 (jim_thompson5910):

1 pound = 16 ounces ... something you memorize or look up 6*(1 pound) = 6*(16 ounces) ... multiply both sides by 6 (since we want 6 pounds) 6 pounds = 96 ounces

OpenStudy (anonymous):

16 oz in a pound so 96

OpenStudy (anonymous):

1 pound = 16 ounces 16 X 6 ANSWER: 96 OUNCES

OpenStudy (mathstudent55):

1 lb = 16 oz 2 * 1 lb = 2 * 16 oz 3 * 1 lb = 3 * 16 oz etc. 6 * 1 lb = 6 * 16 oz 6 lb = 96 oz